A website allows users to download individual songs or an entire album. All individual songs cost the same to download, and all albums cost the same to download. Ryan pays $14.95 to download 5 individual songs and 1 album. Seth pays $22.95 to download 3 individual songs and 2 albums. How much does the website charge to download a song? An entire album?
Answer by Cromlix(4381)

Hi there,
Make an individual song = x
An album = y
Ryan pays $14.95 to download 5 individual songs and 1 album.
5x + y = 14.95........Eq(1)
Seth pays $22.95 to download 3 individual songs and 2 albums.
3x + 2y = 22.95.......Eq(2)
Use simultaneous equations.
5x + y = 14.95........Eq(1)
3x + 2y = 22.95.......Eq(2)
Multiply Eq(1) by 2
10x + 2y = 29.90......Eq(1)
3x + 2y = 22.95.......Eq(2)
Subtract Eq(2) from Eq(1)
7x = 6.95
x = 0.99
Substitute x = 0.99 into Eq (2)
3x + 2y = 22.95.......Eq(2)
3(0.99) + 2y = 22.95
2.98 + 2y = 22.95
2y = 22.95 - 2.98
2y = 19.97
y = 9.99
Individual song = $0.99
Album = $9.99
Hope this helps :-)